Calculate the percentage yield of ethanol when 100 g of glucose is fermented to make 20 g of ethanol.


1
Expert's answer
2021-01-06T05:55:13-0500

the normal reaction of glucose fermentation to ethanol is


"C_6H_{12}O_{6(aq)}\to 2C_2H_5OH_{(aq)} +2CO_{2(g)}"


from the reaction above,

1 mole of glucose ferments to form 2 moles of ethanol

This means that, 180g of glucose ferments to form 92g (2×46) of ethanol


100g of glucose should therefore ferment to form 51.1g ("\dfrac{100×92}{180}") of ethanol



from, "\% yield = \dfrac{\textsf{Actual Yield}}{\textsf{Theoretical Yield}} × 100\%"


"\%yield = \dfrac{20g}{51.1g} × 100\% = 39.14"%



This means that the percentage yield of ethanol when 100 g of glucose is fermented is 39.14%